OpenStudy (anonymous):

what is the derivative of y=2247-1123*sqrt(4-x) This is a distance formula and I am taking the derivative to get a velocity formula

OpenStudy (anonymous):

I would convert the square root into a term with an exponent of 1/2 sqrt(4-x) -> (4-x)^1/2 and use the exponent rule for derivatives

OpenStudy (ranga):

Power rule followed by chain rule.

OpenStudy (anonymous):

what would happen to the 2247?

OpenStudy (anonymous):

the 1123 would become halved whenever you do power rule correct?

OpenStudy (nincompoop):

what is the derivative of a constant? ALWAYS zero this is fundamental and should be part of the definition before knowing all the rules

OpenStudy (nincompoop):

if you doubt me, I can show you the proof that the derivative of a constant is zero or you can check any introduction analysis book like Apostol's Calculus or Spivak Calculus

OpenStudy (anonymous):

that's just odd because that would mean I was getting a negative number

OpenStudy (nincompoop):

f'(x) = -x^2 would be -2x correct?

OpenStudy (anonymous):

right. using power rule

OpenStudy (anonymous):

when i took the derivative i got dy/dt=-561.5(4-x)^-1/2

OpenStudy (nincompoop):

remember that negative exponent gives you \[x^{-1/n}=\frac{ 1 }{ x^{1/n} }\]

OpenStudy (nincompoop):

y'(x) = 1123/(2 sqrt(4-x))

OpenStudy (anonymous):

why isn't the 1123 negative?

OpenStudy (nincompoop):

because there's a negative and a negative \[f'(x) = \frac{ 1 }{ x }=\frac{ -1 }{ x^2 }\]

OpenStudy (anonymous):

ohhhhh ok. I gotchya! I was hung up on that one part. I understand what's going on now. Thanks!!

OpenStudy (nincompoop):

d/dx(y) = d/dx(2247-1123 sqrt(4-x)) The derivative of y is y'(x): y'(x) = d/dx(2247-1123 sqrt(4-x)) Differentiate the sum term by term and factor out constants: y'(x) = d/dx(2247)-1123 d/dx(sqrt(4-x)) The derivative of 2247 is zero: y'(x) = -1123 (d/dx(sqrt(4-x)))+0 Simplify the expression: y'(x) = -1123 (d/dx(sqrt(4-x))) Using the chain rule, d/dx(sqrt(4-x)) = ( dsqrt(u))/( du) ( du)/( dx), where u = 4-x and ( d)/( du)(sqrt(u)) = 1/(2 sqrt(u)): y'(x) = -1123 (d/dx(4-x))/(2 sqrt(4-x)) Differentiate the sum term by term and factor out constants: y'(x) = -(1123 d/dx(4)-d/dx(x))/(2 sqrt(4-x)) The derivative of 4 is zero: y'(x) = -(1123 (-(d/dx(x))+0))/(2 sqrt(4-x)) Simplify the expression: y'(x) = (1123 (d/dx(x)))/(2 sqrt(4-x)) The derivative of x is 1: y'(x) = (1 1123)/(2 sqrt(4-x))

OpenStudy (nincompoop):

even wolframalpha confirms it LOL

OpenStudy (anonymous):

haha well thank you!! now to finish the word problem!